OpenStudy (mislovely187):

What is the equation of a circle with center (5,3) and radius 5?

Vocaloid (vocaloid):

(x-h)^2 + (y - k)^2 = r^2 where h = your x-coordinate, y = your y-coordinate, and r= radius

OpenStudy (welshfella):

I must correct vocaloid her its k = your y-coordinate - just a typo

OpenStudy (welshfella):

just replace h , k and r by the values in your question